RESOLVED: Packet loss through London Internet Exchange (LINX)
Posted by Adam Heavens on 15 December 2010 12:16 PM

Update 12:00 15/12/2010

Following on from the problems experienced yesterday at the London Internet Exchange all issues have now been resolved by LINX and traffic has been restored.

We are continuing to monitor the situation closely and have seen no further problems.

Best regards

Server Centre Support

Read more »

Packet loss through London Internet Exchange (LINX)
Posted by Adam Heavens on 13 December 2010 05:59 PM

Update 10:25 14/12/2010: Following further issues experienced by some of our customers this morning we have taken the action to shutdown all peering with LINX and re-route traffic to our other providers. We will post a further update once LINX have identified the cause of the issues.

Update 22.20: LINX have not reported any further issues therefore we have now re-enabled LINX BGP sessions.

Update 18.00: LINX have informed us that their engineers has fixed the issue, and that normal service has been restored. We have migrated around 80% of our LINX traffic to other providers. We will leave this configuration in place for the next 3 hours to ensure that no repeat issues occur at LINX. Assuming there are no further issues we will move traffic back to LINX at 9pm.


At approximately 16.30 today we were advised by London Internet Exchange (LINX) that they had identified a problem with one of their core switches which is resulting in packet loss. This issue is affecting many UK service providers who peer at LINX.


Poor performance for some routes via LINX.

A number of service providers at LINX have begun to shutdown their BGP sessions at LINX and move traffic to other providers. We intend to shutdown BGP session too.

Server Centre apologises for this issue and are doing everything possible to resolve this issue as quickly as we can.

Further information will be posted as it becomes available.

Best regards,

Server Centre Support

Read more »